# Homological Filling and Minimal Varifolds in Four-Dimensional Einstein Manifolds

### v1

@@ Abstract.
where $F_{\operatorname{Ein}}$ depends only on $(v,D)$ and on quantitative Sobolev and $\varepsilon$–regularity constants for Einstein metrics.
@@ 1. Introduction
This paper is a sequel to our previous work [23], where we considered the larger class with $|{\rm Ric}_{g}|\leq 3,\ \operatorname{Vol}(M,g)\geq v>0,\ \operatorname{diam}(M,g)\leq D,\ H_{1}(M;\mathbb{Z})=0,$ and proved a linear upper bound for the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$. More precisely, for every $(M^{4},g)\in\mathcal{M}_{1}(4,v,D)$ we showed that there exist functions $f_{1}(v,D),f_{2}(v,D)>0$ such that
@@
In the present work we revisit this question in the more rigid Einstein setting. For $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ the curvature solves an elliptic system, and global Sobolev inequalities can be used, following Anderson [2, 3], to derive quantitative $L^{2}$–curvature bounds and $\varepsilon$–regularity with constants depending only on a Sobolev constant. This analytic input, combined with the Einstein specialization of the Cheeger–Naber bubble–tree decomposition and an improved combinatorial filling estimate, allows us to obtain an explicit upper bound for $A_{\min}(M,g)$ in terms of $(v,D)$ and the Sobolev constants.
@@
Moreover $F_{\operatorname{Ein}}(v,D)$ can be written explicitly in terms of the Sobolev constant $C_{S}(v,D)$, the global $L^{2}$–curvature bound, and the quantitative Einstein $\varepsilon$–regularity constants.
@@
Theorem 1.1 is then an immediate consequence of Theorem 1.2 and the min–max theorem of Nabutovsky–Rotman [19], which relates $A_{\min}(M,g)$ to $\operatorname{FH}_{1}$. In the Einstein setting we can track all constants arising from the Sobolev inequality, the $L^{2}$–curvature bound and the bubble–tree decomposition, thus making $F_{\operatorname{Ein}}(v,D)$ completely explicit once the analytic data are fixed.
@@ 2.1. Global Sobolev inequality
(2) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}+\frac{1}{D^{2}}\int_{M}u^{2}\right)^{1/2}.$
@@
Since $|\operatorname{Ric}_{g}|\leq 3$, we have $\operatorname{Ric}_{g}\geq-3g$. By Bishop–Gromov volume comparison, there exists a dimensional constant $c_{0}>0$ such that
$\operatorname{Vol}B_{1}(x)\ \geq\ c_{0}\,v\qquad\text{for all }x\in M.$
In particular $M$ is non-collapsed on the unit scale and has a uniform lower bound on the isoperimetric constant: there exists $I(v,D)>0$ such that for all smooth domains $\Omega\subset M$ with volume at most $\frac{1}{2}\operatorname{Vol}(M)$ one has
$\operatorname{Vol}(\Omega)^{3/4}\ \leq\ I(v,D)\,\mathcal{H}^{3}(\partial\Omega),$
where $\mathcal{H}^{3}$ denotes the $3$-dimensional Hausdorff measure induced by $g$.
By a standard argument (see e.g. Hebey, Nonlinear analysis on manifolds), such inequality is equivalent to an $L^{2}$-Sobolev inequality: there exists $C_{S}(v,D)$ such that
(3) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}\right)^{1/2}$
for every $u\in C^{\infty}(M)$ with zero mean, $\int_{M}u=0$.
For a general $u$ write $u=(u-\bar{u})+\bar{u}$, where $\bar{u}=\operatorname{Vol}(M)^{-1}\int_{M}u$. Then
[…]
The first term is controlled by (3). For the second, we note that
$|\bar{u}|^{2}\leq\frac{1}{\operatorname{Vol}(M)}\int_{M}|u|^{2}\leq\frac{1}{v}\int_{M}|u|^{2},$
and hence
$\Bigl(\int_{M}|\bar{u}|^{4}\Bigr)^{1/2}=|\bar{u}|^{2}\,\operatorname{Vol}(M)^{1/2}\leq\frac{\operatorname{Vol}(M)^{1/2}}{v}\int_{M}|u|^{2}\leq\frac{D}{v}\,\frac{1}{D^{2}}\int_{M}|u|^{2}.$
Absorbing all multiplicative constants into $C_{S}(v,D)$ yields (2). ∎
@@
As in [8], one may obtain the global $L^{2}$ curvature bound for manifolds with our assumptions. In the case of Einstein metric, the proof is simpler due to the Chern–Gauss–Bonnet formula.
Theorem 2.3 ([8], Theorem 1.13).
@@
Since $M$ is Einstein, $\operatorname{Ric}_{g}=\lambda g$ with $|\lambda|\leq 3$, so the scalar curvature is constant and equal to $\mathrm{Scal}=4\lambda$. In dimension $4$ the Riemann tensor decomposes as
[…]
where $W$ is the Weyl tensor and $E$ the traceless Ricci tensor. For Einstein metrics we have $E\equiv 0$, hence
(4) $|\operatorname{Rm}|^{2}=|W|^{2}+\frac{\mathrm{Scal}^{2}}{24}.$
The Chern–Gauss–Bonnet formula in dimension $4$ reads
$8\pi^{2}\chi(M)=\int_{M}\Bigl(|W|^{2}-\frac{1}{2}|E|^{2}+\frac{\mathrm{Scal}^{2}}{24}\Bigr)\,d\operatorname{vol}_{g}.$
Again, $E\equiv 0$ for Einstein metrics, so combining this with (4) gives
$\int_{M}|\operatorname{Rm}|^{2}\,d\operatorname{vol}_{g}=8\pi^{2}\chi(M).$
By the Cheeger–Naber finiteness theorem for Einstein 4-manifolds with two-sided Ricci bounds, a lower volume bound and an upper diameter bound (see [8] together with Anderson [2, 3]), the class $\mathcal{E}_{1}(4,v,D)$ contains only finitely many diffeomorphism types. In particular,
$|\chi(M)|\ \leq\ C_{\chi}(v,D)$
for some constant $C_{\chi}(v,D)$ depending only on $v,D$. It follows that
$\int_{M}|\operatorname{Rm}|^{2}\ =\ 8\pi^{2}|\chi(M)|\ \leq\ 8\pi^{2}C_{\chi}(v,D)$
for all $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$, so we may set $C_{L^{2}}(v,D)=8\pi^{2}C_{\chi}(v,D)$. ∎
@@
2.3. $\varepsilon$-regularity via Sobolev and Moser iteration
We now state and prove a standard Moser-iteration lemma adapted to the inequality (6) and the local Sobolev inequality (8).
Restricting (2) to functions supported in a ball $B_{2r}(x)\subset M$ (with $r\leq D$) immediately gives a local inequality of the form
(7) $\Bigl(\int_{B_{2r}(x)}|f|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{B_{2r}(x)}|\nabla f|^{2}+\frac{1}{r^{2}}\int_{B_{2r}(x)}f^{2}\right)^{1/2}$
for all $f\in C_{c}^{\infty}(B_{2r}(x))$, since $1/D^{2}\leq 1/r^{2}$ when $r\leq D$.
Lemma 2.4 (Local $L^{\infty}$ control from small $L^{2}$).
Let $u\geq 0$ satisfy
$\Delta u\leq C_{0}(u^{2}+u)\quad\text{weakly on }B_{2r}(x)\subset M.$
Assume that on $B_{2r}(x)$ an $L^{2}$-Sobolev inequality of the form
(8) $\Bigl(\int_{B_{2r}(x)}|f|^{4}\Bigr)^{1/2}\leq C_{S}^{\prime}\left(\int_{B_{2r}(x)}|\nabla f|^{2}+\frac{1}{r^{2}}\int_{B_{2r}(x)}f^{2}\right)^{1/2}$
holds for all $f\in C_{c}^{\infty}(B_{2r}(x))$, with $C_{S}^{\prime}$ depending only on $C_{S}(v,D)$. Then there exist constants $\varepsilon_{0}=\varepsilon_{0}(C_{0},C_{S}^{\prime})>0$ and $C_{1}=C_{1}(C_{0},C_{S}^{\prime})>0$ such that if
$\int_{B_{2r}(x)}u^{2}\leq\varepsilon_{0},$
@@ 2.3. $\varepsilon$-regularity via Sobolev and Moser iteration
$\sup_{B_{r}(x)}u\leq C_{1}r^{-2}.$
@@
We give the standard Moser iteration argument. Fix a smooth cutoff function $\eta$ with
$0\leq\eta\leq 1,\quad\eta\equiv 1\text{ on }B_{r}(x),\quad\operatorname{supp}\eta\subset B_{2r}(x),\quad|\nabla\eta|\leq\frac{2}{r}.$
Let $p\geq 2$ and consider the test function $\varphi=\eta^{2}u^{p-1}$ in the weak formulation of (6). Multiplying by $\varphi$ and integrating over $B_{2r}(x)$ gives
$\int\langle\nabla u,\nabla\varphi\rangle\leq C_{0}\int\varphi(u^{2}+u),$
where all integrals are over $B_{2r}(x)$ and we suppress the domain in the notation. Expanding the left-hand side,
$\nabla\varphi=2\eta u^{p-1}\nabla\eta+(p-1)\eta^{2}u^{p-2}\nabla u,$
so
$\int\langle\nabla u,\nabla\varphi\rangle=(p-1)\int\eta^{2}u^{p-2}|\nabla u|^{2}+2\int\eta u^{p-1}\langle\nabla u,\nabla\eta\rangle.$
Using Young’s inequality, for any $\delta>0$,
$2\left|\int\eta u^{p-1}\langle\nabla u,\nabla\eta\rangle\right|\leq\delta\int\eta^{2}u^{p-2}|\nabla u|^{2}+\frac{4}{\delta}\int u^{p}|\nabla\eta|^{2}.$
Choose $\delta=\frac{p-1}{2}$ to absorb half of the main term. Then
$\frac{p-1}{2}\int\eta^{2}u^{p-2}|\nabla u|^{2}\leq C_{0}\int\eta^{2}u^{p+1}+C_{0}\int\eta^{2}u^{p}+\frac{8}{p-1}\int u^{p}|\nabla\eta|^{2}.$
Now set $v=\eta u^{p/2}$. Then
$|\nabla v|\leq\eta u^{p/2-1}|\nabla u|+u^{p/2}|\nabla\eta|,$
and one checks that
$\int|\nabla v|^{2}\leq C(p)\int\eta^{2}u^{p-2}|\nabla u|^{2}+C(p)\int u^{p}|\nabla\eta|^{2}$
for some $C(p)$ depending only on $p$. Combining the two inequalities yields
(9) $\int|\nabla v|^{2}\leq C(p)\Bigl(\int\eta^{2}u^{p+1}+\int\eta^{2}u^{p}+\int u^{p}|\nabla\eta|^{2}\Bigr).$
We now apply the local Sobolev inequality (8) to $f=v$:
$\Bigl(\int|v|^{4}\Bigr)^{1/2}\leq C_{S}^{\prime}\left(\int|\nabla v|^{2}+\frac{1}{r^{2}}\int v^{2}\right)^{1/2}.$
Since $v^{2}=\eta^{2}u^{p}\leq u^{p}$ and $|\nabla\eta|\leq 2/r$, (9) and (8) together give
$\Bigl(\int|v|^{4}\Bigr)^{1/2}\leq C_{2}(p,C_{S}^{\prime})\left(\int u^{p+1}+\int u^{p}+\frac{1}{r^{2}}\int u^{p}\right)^{1/2},$
where $C_{2}$ depends only on $p$ and $C_{S}^{\prime}$. Using the smallness of $\int u^{2}$ and Hölder interpolation, one can absorb the $u^{p+1}$ term. More precisely, for $p\geq 2$ we have
$\int u^{p+1}\leq\|u\|_{L^{2}}^{1-\theta}\|u\|_{L^{q}}^{\theta(p+1)}$
with suitable $q>p$ and $\theta\in(0,1)$. Taking $\varepsilon_{0}>0$ sufficiently small so that $\|u\|_{L^{2}(B_{2r})}\leq\varepsilon_{0}$ makes the $u^{p+1}$ term a small perturbation, which can be absorbed into the left-hand side after a standard iteration. The upshot is that there exist exponents $p_{j}$ with $p_{0}=2$ and $p_{j+1}=2p_{j}$ and radii $r_{j}$ with $r\leq r_{j}\leq 2r$, and a constant $C_{3}=C_{3}(C_{0},C_{S}^{\prime})$ such that
$\|u\|_{L^{p_{j+1}}(B_{r_{j+1}})}\leq C_{3}\,r^{-2}\,\|u\|_{L^{p_{j}}(B_{r_{j}})}\quad\text{for all }j\geq 0.$
Iterating this inequality and letting $j\to\infty$ yields
$\|u\|_{L^{\infty}(B_{r}(x))}\leq C_{1}(C_{0},C_{S}^{\prime})\,r^{-2}\,\|u\|_{L^{2}(B_{2r}(x))}.$
If we now assume $\int_{B_{2r}(x)}u^{2}\leq\varepsilon_{0}$ with $\varepsilon_{0}$ fixed, we obtain
$\sup_{B_{r}(x)}u\leq C_{1}r^{-2},$
as claimed. ∎
Combining this with Theorem 2.1 (restricted to $B_{2r}(x)$ as in (7)) and the curvature equation we obtain:
Theorem 2.5 (Einstein $\varepsilon$-regularity via Sobolev).
There exist dimensional constants $\varepsilon_{\mathrm{Ein}}>0$ and $c_{\mathrm{Ein}}>0$, depending only on $C_{S}(v,D)$, such that the following holds.
$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}.$
Then
(10) $\sup_{B_{r}(x)}|\operatorname{Rm}|\leq Cr^{-2}$
for some universal constant $C$, and the harmonic radius at every point $y\in B_{r}(x)$ satisfies
$r_{h}(y)\geq c_{\mathrm{Ein}}\,r.$
Proof.
Set $u=|\operatorname{Rm}|$. By (6) we have $\Delta u\leq C_{0}(u^{2}+u)$ on $B_{2r}(x)$. The restriction of the global Sobolev inequality (2) to $B_{2r}(x)$ yields a local inequality of the form (8) with $C_{S}^{\prime}=C_{S}(v,D)$, as in (7).
Applying Lemma 2.4, we obtain constants $\varepsilon_{\mathrm{Ein}}=\varepsilon_{0}(C_{0},C_{S}(v,D))>0$ and $C_{1}=C_{1}(C_{0},C_{S}(v,D))$ such that
$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}\quad\Longrightarrow\quad\sup_{B_{r}(x)}|\operatorname{Rm}|\leq C_{1}r^{-2},$
which is (10).
To obtain the harmonic radius bound, we use standard elliptic regularity in harmonic coordinates. Fix $y\in B_{r}(x)$. Choose a harmonic coordinate chart $\Phi:B_{\rho}(0)\subset\mathbb{R}^{4}\to M$ with $\Phi(0)=y$, where $\rho>0$ is chosen so small that $B_{2\rho}(y)\subset B_{r}(x)$. In harmonic coordinates, the metric components $g_{ij}$ satisfy a quasi-linear elliptic system of the form
$\Delta g_{ij}=Q_{ij}(g,\partial g)-2\operatorname{Ric}_{ij},$
where $Q$ is quadratic in $\partial g$ and $\operatorname{Ric}_{ij}=\lambda g_{ij}$ is bounded. The curvature bound (10) on $B_{r}(x)$ in turn bounds the coefficients of this system. By standard Schauder or $W^{2,p}$ interior estimates (see, e.g., Anderson [2, 3]), one obtains a uniform $C^{1,\alpha}$ bound on $g_{ij}$ in $B_{\rho/2}(0)$, with constants depending only on the curvature bound and $C_{S}(v,D)$. In particular, by rescaling we can choose $\rho$ comparable to $r$ so that
$\|g_{ij}-\delta_{ij}\|_{C^{0}(B_{\rho/2}(0))}+\rho\,\|\partial g_{ij}\|_{C^{0}(B_{\rho/2}(0))}\leq 10^{-3}.$
This implies that the harmonic radius at $y$ satisfies $r_{h}(y)\geq c_{\mathrm{Ein}}\,r$ for some $c_{\mathrm{Ein}}>0$ depending only on $C_{S}(v,D)$ and the dimension. ∎
Corollary 2.6 (Harmonic radius scale).
Define the dimensionless parameter
$\mathbf{A}(v,D):=\frac{C_{L^{2}}(v,D)}{\varepsilon_{\mathrm{Ein}}},$
where $C_{L^{2}}(v,D)$ is as in Theorem 2.3 and $\varepsilon_{\mathrm{Ein}}$ is as in Theorem 2.5. There exists a decreasing function $\Psi:(0,\infty)\to(0,\infty)$ depending only on $C_{S}(v,D)$ and the dimension, such that for each $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ one can cover $M$ by finitely many balls $B_{r_{i}}(x_{i})$ with
$\int_{B_{2r_{i}}(x_{i})}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}},$
and hence
$r_{h}(y)\geq r_{0}(v,D)\,r_{i}\quad\text{for all }y\in B_{r_{i}}(x_{i}),\qquad r_{0}(v,D):=\Psi(\mathbf{A}(v,D)).$
Proof.
Fix $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$. For each $x\in M$, define
$r(x):=\sup\Bigl\{\,0<\rho\leq D:\ \int_{B_{2\rho}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}\,\Bigr\}.$
By continuity of the map $\rho\mapsto\int_{B_{2\rho}(x)}|\operatorname{Rm}|^{2}$ and compactness of $M$, we have $r(x)>0$ for all $x$ and the family $\{B_{r(x)}(x)\}_{x\in M}$ is an open cover of $M$. By compactness, there is a finite subcover $\{B_{r_{i}}(x_{i})\}_{i=1}^{N}$.
On each ball $B_{2r_{i}}(x_{i})$ we have, by construction,
$\int_{B_{2r_{i}}(x_{i})}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}.$
Thus Theorem 2.5 applies (after possibly replacing $r_{i}$ by $r_{i}/2$), and we obtain
$r_{h}(y)\geq c_{\mathrm{Ein}}\,r_{i}\qquad\text{for all }y\in B_{r_{i}}(x_{i}).$
We may now define $\Psi$ by setting
$\Psi(\mathbf{A}(v,D)):=c_{\mathrm{Ein}},$
viewed as a function of the pair $(C_{S}(v,D),\varepsilon_{\mathrm{Ein}})$ and hence of $\mathbf{A}(v,D)=C_{L^{2}}(v,D)/\varepsilon_{\mathrm{Ein}}$. By enlarging $\Psi$ if necessary for small values of $\mathbf{A}$, we can assume that $\Psi$ is decreasing on $(0,\infty)$. Setting $r_{0}(v,D)=\Psi(\mathbf{A}(v,D))$ yields the claimed statement. ∎
@@
In all subsequent estimates we write $r_{0}(v,D)$ for this harmonic-radius scale. Ultimately, all constants will be expressed as functions of $C_{S}(v,D)$ and $\mathbf{A}(v,D)$.

### v2

@@ Abstract.
where $F_{\operatorname{Ein}}$ depends only on $(v,D)$. The proof uses the Einstein-specific analytic input of a uniform Sobolev inequality, a global $L^{2}$-curvature bound, and standard $\varepsilon$-regularity/harmonic-radius estimates.
@@ 1. Introduction
This paper is a sequel to our previous work [24], where we considered the larger class with $|{\rm Ric}_{g}|\leq 3,\ \operatorname{Vol}(M,g)\geq v>0,\operatorname{diam}(M,g)\leq D,\ H_{1}(M;\mathbb{Z})=0,$ and proved an affine upper bound for the first homological filling function $\operatorname{FH}_{1}$ of $(M,g)$. More precisely, for every $(M^{4},g)\in\mathcal{M}_{1}(4,v,D)$ we showed that there exist functions $f_{1}(v,D),f_{2}(v,D)>0$ such that
@@
In the present work we revisit this question in the more rigid Einstein setting. For $(M^{4},g)\in\mathcal{E}_{1}(4,v,D)$ the curvature solves an elliptic system, and global Sobolev inequalities can be used, following Anderson [2, 3], to derive quantitative $L^{2}$–curvature bounds and $\varepsilon$–regularity with constants depending only on a Sobolev constant. This analytic input, combined with the Einstein specialization of the Cheeger–Naber bubble–tree decomposition and an improved combinatorial filling estimate, allows us to obtain a quantitative upper bound for $A_{\min}(M,g)$ in terms of $(v,D)$ and the Sobolev constants.
@@
Moreover, the dependence of $F_{\operatorname{Ein}}(v,D)$ is reduced to the Sobolev constant $C_{S}(v,D)$, the global $L^{2}$-curvature bound, and the standard Einstein $\varepsilon$-regularity constants.
@@
Theorem 1.1 is then an immediate consequence of Theorem 1.2 and the min–max theorem of Nabutovsky–Rotman [20], which relates $A_{\min}(M,g)$ to $\operatorname{FH}_{1}$. In the Einstein setting we can track the dependence of the constants arising from the Sobolev inequality, the $L^{2}$-curvature bound and the bubble-tree decomposition, thus reducing the dependence of $F_{\operatorname{Ein}}(v,D)$ to these analytic and combinatorial inputs.
@@ 2.1. Global Sobolev inequality
(2) $\Bigl(\int_{M}|u|^{4}\Bigr)^{1/2}\ \leq\ C_{S}(v,D)\left(\int_{M}|\nabla u|^{2}+\frac{1}{D^{2}}\int_{M}u^{2}\right).$
@@
Under our standing assumptions we have $\operatorname{Ric}_{g}\geq-3g$, $\operatorname{Vol}(M)\geq v$ and $\operatorname{diam}(M)\leq D$. A standard consequence is a uniform isoperimetric inequality, hence a uniform Sobolev inequality for mean-zero functions; see, for instance, [16, 2]. Thus there exists $C^{\prime}_{S}(v,D)$ such that
$\|u-\bar{u}\|_{L^{4}(M)}^{2}\leq C^{\prime}_{S}(v,D)\,\|\nabla u\|_{L^{2}(M)}^{2},\qquad\bar{u}:=\operatorname{Vol}(M)^{-1}\int_{M}u.$
In addition, Bishop–Gromov comparison with $\operatorname{Ric}_{g}\geq-3g$ and $\operatorname{diam}(M)\leq D$ gives an upper volume bound $\operatorname{Vol}(M)\leq V_{+}(D)$. Since
$|\bar{u}|^{2}\leq\operatorname{Vol}(M)^{-1}\|u\|_{L^{2}(M)}^{2}\leq v^{-1}\|u\|_{L^{2}(M)}^{2},$
we obtain
$\|\bar{u}\|_{L^{4}(M)}^{2}=|\bar{u}|^{2}\,\operatorname{Vol}(M)^{1/2}\leq v^{-1}V_{+}(D)^{1/2}\,\|u\|_{L^{2}(M)}^{2}\leq C^{\prime\prime}_{S}(v,D)\,\frac{1}{D^{2}}\|u\|_{L^{2}(M)}^{2},$
after enlarging the constant. Combining the bounds for $u-\bar{u}$ and $\bar{u}$ yields (2). ∎
@@
We use the local $L^{2}$ curvature estimate of Cheeger–Naber together with a covering argument to obtain a global bound in our setting.
Theorem 2.3 ($L^{2}$ curvature bound).
@@
Since $|\operatorname{Ric}_{g}|\leq 3$, Bishop–Gromov comparison and the bounds $\operatorname{Vol}(M)\geq v$, $\operatorname{diam}(M)\leq D$ imply that there exists $v_{1}(v,D)>0$ such that
$\operatorname{Vol}(B_{1}(x))\geq v_{1}(v,D)\qquad\text{for every }x\in M.$
Applying the local $L^{2}$ estimate of Cheeger–Naber [8, Theorem 1.13] to each unit ball gives
$\int_{B_{1}(x)}|\operatorname{Rm}|^{2}\leq C_{\mathrm{CN}}(v_{1}(v,D))\qquad\text{for every }x\in M.$
Now choose a maximal $\frac{1}{2}$-separated set $\{x_{1},\dots,x_{N}\}\subset M$. Then the balls $B_{1}(x_{i})$ cover $M$, while the balls $B_{1/4}(x_{i})$ are disjoint. Using again Bishop–Gromov comparison together with the diameter bound, we obtain $N\leq N(v,D)$ for some constant depending only on $(v,D)$. Therefore
$\int_{M}|\operatorname{Rm}|^{2}\leq\sum_{i=1}^{N}\int_{B_{1}(x_{i})}|\operatorname{Rm}|^{2}\leq N(v,D)\,C_{\mathrm{CN}}(v_{1}(v,D)).$
This gives the desired global bound. ∎
@@
2.3. Einstein $\varepsilon$-regularity
The next result is standard in dimension four; we record it in the form needed later.
Theorem 2.4 (Einstein $\varepsilon$-regularity).
There exist constants $\varepsilon_{\mathrm{Ein}}(v,D)>0$, $C_{\mathrm{Ein}}(v,D)>0$ and $c_{\mathrm{Ein}}(v,D)>0$ such that the following holds.
$\int_{B_{2r}(x)}|\operatorname{Rm}|^{2}\leq\varepsilon_{\mathrm{Ein}}(v,D).$
@@ 2.3. Einstein $\varepsilon$-regularity
$\sup_{B_{r}(x)}|\operatorname{Rm}|\leq C_{\mathrm{Ein}}(v,D)\,r^{-2},\qquad r_{h}(y)\geq c_{\mathrm{Ein}}(v,D)\,r\quad\text{for all }y\in B_{r}(x).$
@@
After rescaling we may assume $r=1$. The hypotheses place us in the standard noncollapsed Einstein setting: the Ricci tensor is uniformly bounded, local volume ratios are bounded below in terms of $(v,D)$, and the Sobolev constant is controlled by Theorem 2.1. The usual $\varepsilon$-regularity theorem for Einstein $4$-manifolds therefore applies and yields the pointwise curvature bound together with the harmonic-radius estimate on $B_{1}(x)$; see Anderson [2, 3], Cheeger–Tian [9], and Cheeger–Naber [8]. Rescaling back proves the statement. ∎
@@
In what follows we only use the existence of the constants in Theorems 2.1, 2.3 and 2.4. The global harmonic-radius scale entering the bubble–tree decomposition will be taken directly from Cheeger–Naber’s theorem.